# HoneySat: A Network-based Satellite Honeypot Framework

### v1

@@ 6.4 Internet Interaction Experiment
Honeypot 1 Egypt 4 2 hr
Honeypot 2 Sweden (Tor) 9 5 min
Honeypot 1 France (Tor) 6 4 min
Honeypot 3 USA 8 3 min
Table 5: Exposed Telnet interactions received.
Experiment Results. Our honeypot servers were successfully indexed by Shodan. Both Telnet and web banners were captured, and none of them were tagged as honeypots.
Our honeypots successfully enticed external actors and captured 4 distinct sessions (show in Table 5) via the exposed Telnet protocol discussed in Sec. 4.3. The exposed Telnet protocol was implemented so that a human had to type “activate” in the terminal before sending commands. This feature was implemented to filter out Internet Telnet bots. As such, the commands that we describe below were sent by a human and not a crawler or Internet bot.
We now describe the commands received in the first interaction session shown in Table 5.
  1. 1.
    help: Shows the ground software available commands.
  2. 2.
    ls: The Linux ls command is not a valid command in the honeypot ground software.
  3. 3.
    fp_show: Prints the list of commands in the current flight plan (The flight plan refers to a pre-programmed sequence of tasks, and commands that guide the satellite’s operation).
  4. 4.
    com_debug: Prints the current CSP network configuration including the current CSP node, interfaces and routing table.
The commands fp_show and com_debug are flight software-specific commands which suggests that the adversary was successfully deceived by our honeypot and purposefully sent flight software commands instead of haphazardly sending unrelated commands. Another factor that indicates that the adversary was deceived is that they engaged with our honeypot over a period of two hours indicating that they did not immediately identified it as a honeypot.
@@
Key finding Q-2 Human adversaries interacted with three of our HoneySat-powered honeypots during four distinct Telnet sessions resulting in 16 satellite mission-specific commands captured.
6.5 Case Study: Generic CCSDS Mission Honeypot
In Sec. 5.2 we described how we implemented HoneySat using one SmallSat protocol ecosystem, namely, CSP. In this case study we are interested in testing the extensibility capabilities of HoneySat to support additional ecosystems (discussed in Sec. 2.5) by adding a second ecosystem to HoneySat, namely the CCSDS ecosystem.
@@ 6.5 Case Study: Generic CCSDS Mission Honeypot
Methodology. Building upon our HoneySat framework implementation, as detailed in Sec. 5, we enhanced the system by integrating various components of the RACCOON OS and the YAMCS framework.
Results. We successfully extended HoneySat to support the CCSDS ecosystem. Regarding the ground segment we implemented the exposed network protocol using YAMCS’ built-in web interface (Fig. 7), for the mission control software we used YAMCS’s built-in Mission Control Software [1], for the radio simulator we used RACCOON’s communication application and for the ground configuration and logging repository we again used YAMCS built-in features.
Focusing on the space segment, we implemented the satellite flight software runtime using the RACCOON framework, and for the Flight Software Services, we configured the RACCOON framework to connect it to the YAMCS’s MCS on the ground segment. The satellite personality and logging repository were based on the existing HoneySat implementations.
The majority of the effort involved in extending HoneySat to support the CCSDS ecosystem was dedicated to understanding the ecosystem itself, including components such as PUS. Additionally, we had to analyze the RACCOON flight software code and the YAMCS framework documentation. The only extra implementation that was required was the Flight Software Services which we modified using Rust. Other than that we reused several elements of HoneySat like the satellite personality and the Satellite Simulator.
@@
7 Discussion and Future Work
@@
Broader Applications of HoneySat. While originally designed as a honeypot, our framework offers the flexibility to support a range of applications beyond its initial purpose. One promising use case is the development of digital twins for satellite systems, enabling the simulation of real-world satellite subsystems and communication scenarios. Furthermore, HoneySat can be integrated into cyber range environments to enhance training programs and cybersecurity simulation exercises. Lastly, HoneySat serves as an educational tool, providing researchers, industry professionals, and security enthusiasts with an opportunity to explore and learn about satellite architecture and operations.
8 Conclusion
Although we have yet to witness a Stuxnet-like cyberattack on space systems, security researchers and professionals need to develop effective countermeasures to secure satellites. In this paper, we introduced HoneySat, the first satellite honeypot that provides a much-needed alternative source of empirical data on attackers’ TTPs. We created a satellite honeypot that realistically simulates an entire satellite system, including its sensors and subsystems.
We performed several experiments that provide strong evidence that the framework can obtain rich interaction data, can be extended and customized, and simulate a satellite to keep its honeypot nature hidden from potential adversaries. Finally, we hope that security researchers and professionals take advantage of HoneySat’s open-source implementation and use it as a foundation not only for satellite honeypot deployments but also for simulation, education, and training applications.
9 Ethical Considerations
In this paper we consider the ethical consequences and possible negative outcomes of two main elements of our research. The satellite operator user study and the multiple deployments of our satellite honeypot exposed to the Internet. We now discuss the stakeholders, potential risks and how we mitigated those risks for each of these elements.
9.1 Satellite Operators User Study
Our user study adheres to the ethical guidelines outlined in the USENIX Security ‘25 Ethics Policy. Prior to conducting any research, the study plan, survey protocol, and associated materials were submitted for review by our Institutional Review Board (IRB), which evaluated the study’s potential risks to participants. The protocol received "exempt status" from the IRB, indicating that the research involves no more than minimal risk. This status reflects our careful evaluation of possible harms, which we discuss further below.
@@ 9.1 Satellite Operators User Study
Risk Mitigation.The main risk for the survey participants is the breach of privacy. To mitigate this risk, we ensured that no identifiable data (e.g., names, contact details) was collected from survey participants and the survey responses remain fully anonymous. Additionally, the survey itself includes an informed consent section that informs the respondents about the voluntary participation, the survey’s purpose, a description of the procedures, the risks involved, and contact information and the option to opt out of the survey. Finally, we provided participants the choice to skip questions using the response options “Prefer not to say” and “I do not know” to applicable questions. Due to the aforementioned risk mitigation strategies the stakeholders’ breach of privacy risk is negligible.
9.2 Honeypot Deployment Experiment
@@
Risk Mitigation. The main risk for the infrastructure owners is that one of the external actors will use our honeypot as an stepping stone to breach their production infrastructure. For the deployment on the premises of the institutions we worked with the local IT infrastructure administrators to ensure that we took all the necessary precautions to avoid any possibility that the external actors could gain access to the infrastructure via our honeypot. First, the infrastructure owners provided us with a server with a clean OS installation that did not have any production applications or sensitive data. Second, the server was isolated on its own network segment. Third, we configured a firewall to only allow the necessary ports. Fourth, the VNC portion of the honeypot is configured so that external actors cannot use it to host malicious services by denying traffic forwarding. Fifth and final, we deployed our honeypot using two sandboxing layers (Docker containers and Virtual Machines (VMs) as we discussed in Appendix A.3. Due to the aforementioned risk mitigation strategies the stakeholders’ breach of privacy risk is negligible.
The main risk for the external actors is the breach of their privacy. However, it is generally accepted that trespassers of a computer system do not have reasonable expectation of privacy [66]. In order to mitigate this risk our honeypot system gives notice and warning to any user connecting to it indicating that "This computer system is for authorized use only." Additionally, the web services of our honeypot are protected by a username and password. This notice system gives external actors a chance to avoid interaction, which helps reduce the risk.
@@ Appendix A Appendix
A.1 Interaction Sequences and Exploits
Table 6 includes all the interaction sequences and exploits we performed during the experiments in Sec. 6.2.
Table 6: Tactics, Techniques and Procedures’ Interaction Experimental Exploits.
@@ A.1 Interaction Sequences and Exploits
A.2 Satellite Simulator Design Details
@@
These simulation modules can communicate with each other to facilitate proper functionality. For example, the power simulation module will query the orbital simulation for orbital data to determine if the satellite is positioned properly in relation to the sun so that it can draw power from its solar panels. The subsystems and the sensors of the Satellite Simulator are designed around these simulations. The Satellite Simulator includes the orbital, rotation, power system, thermal, magnetic, and payload simulations. Due to space restrictions and their complex design, we explain each simulation in Appendix A.2.
@@
Subsystem State Modules. These modules are abstractions of the satellite subsystems discussed in Sec. 2.4. A given subsystem state is made up of one or more sensors. For example, in Fig. 5 the Payload State (light green) includes the Camera Sensor (light blue). In this way, subsystem states serve two purposes. They can get data from their sensors or can set a specific configuration on their sensors.
@@
  • RGB Camera Simulation. This simulation replicates the functionality of an RGB camera pointed at Earth. It uses Earth observation satellite imagery from the EarthExplorer database from the U.S. Geological Survey (USGS) [76]. If a capture image command is issued, it will select an image from one of two pools of images: daytime or nighttime.
    This simulation communicates with the orbital simulation to determine if satellite is in the presence of the sun.
A.3 Deployment and Security Hardening Implementation
@@
Figure 7: Screenshot of the YAMCS web interface, displaying a TM packet, returned from the Generic CCSDS Mission Honeypot.
As we mentioned in Sec. 2.1, high-interaction honeypots such as HoneySat present a high risk of adversary takeover. To mitigate this risk, we implemented HoneySat with two sandboxing layers.

### v2

@@ VI-D Internet Interaction Experiment
ACS3 United States Egypt 3 2 hr
PIXL-1 Germany Sweden (Tor Node) 6 5 min
ACS3 United States France (Tor Node) 5 4 min
ACS3 United States USA 8 3 min
Experiment Results. Three of our honeypots successfully enticed external actors and captured four distinct interaction sessions (show in Table VI) via the exposed Telnet protocol discussed in Sec. IV-C. The four interaction sessions yielded 22 satellite flight software-specific commands. These flight software-specific commands show that the adversaries that interacted with HoneySat were purposefully trying to exploit our honeypot using commands that the flight software recognized, not just haphazardly sending random commands.
We now describe one of the most complex interaction sessions HoneySat captured and analyze it in terms of the SPACE-SHIELD matrix (discussed in Sec. II-F). In this session, the adversary targeted one of the ACS3 honeypots located in the United States. First, they connected to the exposed Telnet server thus getting access to the ground segment, this corresponds to the Ground Segment Compromise (ID T1584.001) technique. Then they issued the “help” and “test” commands which list the available commands supported by the MCS and flight software; these commands correspond to the Gather Victim Mission Information (ID T2002) technique. Later, the adversary issued the “tm dump 0x0000”, which writes telemetry data to a file, showing an attempt to extract mission data, which matches the Exfiltration Over TM Channel (ID T2022) technique. Next, they sent the “com_ping” command which attempts to initiate an interaction with the target spacecraft, matching the Active Scanning (RF/Optical) (ID T2001) technique. The adversary again implemented the Exfiltration Over TM Channel (ID T2022) technique by issuing the “tm_parse_beacon” command which prints beacon telemetry (lightweight periodic telemetry) and the “obc_get_sensor” command which instructs the OBC to print a particular sensor’s data. Finally, the adversary attempted to tamper with the spacecraft’s OBC by issuing the “obc_update_status” command. This command updates the OBC status variables to any arbitrary value, thus matching the Modification of On Board Control Procedures (ID T2010) technique.
Overall, this interaction session alone involved a chain of five SPACE-SHIELD matrix techniques. However, after analyzing all four interaction sessions, we identified the Spacecraft’s Components Discovery (ID T2034) technique, totaling six SPACE-SHIELD techniques based on real-world data captured by HoneySat after successfully enticing and deceiving human adversaries.
It is worth noting that satellites and space systems in general are far less common compared to CPS such as ICS and UAVs making these interactions collected by HoneySat all the more valuable. Additionally, these interactions were produced by human adversaries and not crawlers. We know this because HoneySat’s exposed Telnet was implemented so that a human had to type “activate” before sending commands effectively filtering out Internet Telnet bots.
@@
Key findings Q-2 HoneySat deceived human adversaries who sent 22 satellite flight software-specific commands. The interactions collected by our HoneySat Internet deployments account for six SPACE-SHIELD matrix techniques.
VI-E Case Study: Extending HoneySat to CCSDS Ecosystem
TABLE VII: SmallSat ecosystems (2) and example SmallSats (3) integrated to HoneySat.
Honeypot
Framework
Generic
SmallSat
Ground
Configuration
Institution
HoneySat CSP
ACS3
NASA
PIXL-1
DLR
SUCHAI-2
U of Chile
CCSDS
PUS
N/A
In Sec. V-B we described how we implemented HoneySat using one SmallSat protocol ecosystem, namely, CSP. In this case study, we are interested in testing the extensibility capabilities of HoneySat to support additional ecosystems (discussed in Sec. II-E) by adding a second ecosystem to HoneySat, namely the CCSDS ecosystem.
@@ VI-E Case Study: Extending HoneySat to CCSDS Ecosystem
Experiment Methodology. Building upon our HoneySat framework implementation, as detailed in Sec. V, we enhanced the system by integrating various components of the RACCOON OS and the YAMCS framework.
Regarding the ground segment, we implemented the exposed network protocol using YAMCS’ built-in web interface (Fig. 8). For the mission control software, we used YAMCS’s built-in Mission Control Software [75]. For the radio simulator, we used RACCOON’s communication application. For both the ground configuration and the logging repository, we again used YAMCS built-in features.
For the space segment, we implemented the flight software runtime using the RACCOON framework. For the flight software services, we configured the RACCOON framework to connect it to the YAMCS’s MCS on the ground segment. The satellite personality and logging repository were based on the existing HoneySat implementations.
Experiment Results. We successfully extended HoneySat to support the CCSDS ecosystem. The implementation was completed by a graduate student with no prior satellite-related in the span of two weeks. However, the majority of the effort involved in extending HoneySat to support the CCSDS ecosystem was dedicated to understanding the ecosystem itself, and analyzing the RACCOON flight software code and the YAMCS framework documentation. The only extra implementation that was required was the flight software services which we modified using Rust. Other than that, we reused several HoneySat modules such as the satellite personality and the Satellite Simulator.
@@
VI-F Case Study: Hardware-in-the-loop Experiment
This experiment is designed to produce evidence of the robustness of HoneySat to support real-world satellite hardware-in-the-loop (HIL) operations.
Experiment Description. In this experiment we integrated HoneySat with an in-orbit, operational SmallSat mission. To achieve this, we collaborated with Quantum Galactics GmbH [76], an aerospace company that develops hardware subsystems for SmallSats. The experiment involved sending a telecommand from HoneySat’s ground segment simulation which was then sent to the in-orbit SmallSat, which would then process the telecommand and respond with sanitized telemetry. Due to confidentiality and security restrictions, we are unable to disclose the name of the in-orbit SmallSat which we refer to as SmallSat X.
Experiment Methodology. In order to achieve the HIL integration, we customized both HoneySat and SmallSat X’s mission. For HoneySat, we modified the Satellite Simulator to support SmallSat X’s mission control software (YAMCS). Specifically, we deployed a proxy server that connected the Satellite Simulator to the SmallSat X’s mission production YAMCS instance. Conversely, the SmallSat X’s environment was customized by Quantum Galactics’ team by deploying a script in their YAMCS instance to complete the integration. Once the integration was completed, we coordinated with Quantum Galactics to send a telecommand from HoneySat during one of SmallSat X’s passes.
Experiment Results. The telecommand successfully reached SmallSat X during one of the passes and HoneySat received the appropriate telemetry. Due to security restrictions, we are not able to disclose further details on the telecommand and the telemetry received. However, the entire communication was initiated by HoneySat which in turn received sanitized telemetry from SmallSat X, effectively closing the loop in the HIL experiment. Additionally, Quantum Galactics provided us with a snippet of decoded radio signal data shown in Fig. 5, which confirms that the experiment was successful. In summary, these results highlight HoneySat’s robust extensibility features and provide strong evidence for answering question Q-3 in the affirmative and design objective DO-3 as achieved.
Figure 5: Waterfall plot that shows the radio signal spectrogram when HoneySat communicated with SmallSat X. The signal on top (blue box) belongs to the telecommand sent by HoneySat, and the signal on the bottom (red box) belongs to the telemetry received from the SmallSat X.
Key findings Q-3 HoneySat was successfully integrated into a satellite hardware-in-the-loop operation and communicated with an in-orbit, operational SmallSat.
VII Discussion and Future Work
HoneySat’s Generalizability. HoneySat was designed to be extensible and modular as we discussed in Sec. IV-A which makes our framework generalizable to additional SmallSat missions. We showed this by extending HoneySat to support the two most widely used SmallSat protocol ecosystems, namely CSP and CCSDS (Sec. VI-E), two flight software, SUCHAI and RACOON (Sec. VI-E), and we customized HoneySat to simulate three different SmallSats (Sec. VI-D).
Additionally, we provide a plethora of customization tools including the Satellite Simulator (Sec IV-D), the satellite personality configuration, the ground configuration, and multiple exposed network protocols (Sec V-B2).
Finally, in Table IX, we provide examples of different extensibility use cases, including high-level steps and the HoneySat modules involved so that researchers can use and extend HoneySat for their own experiments.
@@ VII Discussion and Future Work
Broader Applications of HoneySat. Originally designed as a honeypot, our framework can potentially support a range of applications beyond its initial purpose. One promising use case is the development of digital twins for satellite systems, enabling the simulation of real-world satellite subsystems and communication scenarios. Furthermore, HoneySat can be integrated into cyber range environments to enhance cybersecurity training exercises.
VIII Conclusion
Although we have yet to witness a Stuxnet-like cyberattack on space systems, security researchers need to develop effective countermeasures to secure satellites. In this paper, we introduced HoneySat, the first satellite honeypot that realistically simulates an entire satellite mission and provided strong evidence that shows that our honeypot can obtain real-world interaction data and can be extended and customized with real-world satellite software and hardware. Finally, we hope that security researchers use HoneySat’s open-source implementation as a foundation not only for satellite honeypot deployments but also for simulation, education, and training applications.
IX Ethical Considerations
In this paper, we consider the ethical consequences and possible negative outcomes of two main elements of our research. The satellite operator user study and the multiple deployments of our satellite honeypot exposed to the Internet. We now discuss the stakeholders, potential risks, and how we mitigated those risks for each of these elements.
IX-A Satellite Operators User Study
Prior to conducting any research, the study plan, survey protocol, and associated materials were submitted for review by our Institutional Review Board (IRB), which evaluated the study’s potential risks to participants. The protocol received ”exempt status” from the IRB, indicating that the research involves no more than minimal risk. This status reflects our careful evaluation of possible harms, which we discuss further below.
@@ IX-A Satellite Operators User Study
Risk Mitigation.The main risk for the survey participants is the breach of privacy. To mitigate this risk, we ensured that no identifiable data (e.g., names, contact details) was collected from survey participants, and the survey responses remain fully anonymous. Additionally, the survey itself includes an informed consent section that informs the respondents about the voluntary participation, the survey’s purpose, a description of the procedures, the risks involved, contact information, and the option to opt out of the survey. Finally, we provided participants the choice to skip questions using the response options “Prefer not to say” and “I do not know” to applicable questions. Due to the aforementioned risk mitigation strategies, the stakeholders’ breach of privacy risk is negligible.
IX-B Honeypot Deployment Experiment
@@
Risk Mitigation. The main risk for the infrastructure owners is that one of the external actors will use our honeypot as an stepping stone to breach their production infrastructure. For the deployment on the premises of the institutions, we worked with the local IT infrastructure administrators to ensure that we took all the necessary precautions to avoid any possibility that the external actors could gain access to the infrastructure via our honeypot. First, the infrastructure owners provided us with a server with a clean OS installation that did not have any production applications or sensitive data. Second, the server was isolated on its own network segment. Third, we configured a firewall to only allow the necessary ports. Fourth, the VNC portion of the honeypot is configured so that external actors cannot use it to host malicious services by denying traffic forwarding. Fifth and final, we deployed our honeypot using two sandboxing layers (Docker containers and Virtual Machines (VMs) as we discussed in Appendix A-C. Due to the aforementioned risk mitigation strategies, the stakeholders’ breach of privacy risk is negligible.
The main risk for the external actors is the breach of their privacy. However, it is generally accepted that trespassers of a computer system do not have reasonable expectation of privacy [78]. In order to mitigate this risk, our honeypot system gives notice and warning to any user connecting to it indicating that ”This computer system is for authorized use only.” Additionally, the web services of our honeypot are protected by a username and password. This notice system gives external actors a chance to avoid interaction, which helps reduce the risk.
@@ Appendix A Appendix
A-A Interaction Sequences and Exploits
Table VIII includes all the interaction sequences and exploits we performed during the experiments in Sec. VI-B.
TABLE VIII: Tactics, Techniques and Procedures’ Interaction Experimental Exploits.
@@ A-A Interaction Sequences and Exploits
A-B Satellite Simulator Design Details
@@
These simulation modules can communicate with each other to facilitate proper functionality. For example, the power simulation module will query the orbital simulation for orbital data to determine if the satellite is positioned properly in relation to the sun so that it can draw power from its solar panels. The subsystems and the sensors of the Satellite Simulator are designed around these simulations. The Satellite Simulator includes the orbital, rotation, power system, thermal, magnetic, and payload simulations. Due to space restrictions and their complex design, we explain each simulation in Appendix A-B.
@@
Subsystem State Modules. These modules are abstractions of the satellite subsystems discussed in Sec. II-D. A given subsystem state is made up of one or more sensors. For example, in Fig. 6 the Payload State (light green) includes the Camera Sensor (light blue). In this way, subsystem states serve two purposes. They can get data from their sensors or can set a specific configuration on their sensors.
@@
  • RGB Camera Simulation. This simulation replicates the functionality of an RGB camera pointed at Earth. It uses Earth observation satellite imagery from the EarthExplorer database from the U.S. Geological Survey (USGS) [77]. If a capture image command is issued, it will select an image from one of two pools of images: daytime or nighttime. This simulation communicates with the orbital simulation to determine if satellite is in the presence of the sun.
A-C Deployment and Security Hardening Implementation
@@
As we mentioned in Sec. II-A, high-interaction honeypots such as HoneySat present a high risk of adversary takeover. To mitigate this risk, we implemented HoneySat with two sandboxing layers.
@@
Figure 8: Screenshot of the HoneySat YAMCS web interface, displaying a TM packet, returned from the Generic CCSDS Mission Honeypot.
A-D Extensibility Example Use Cases
TABLE IX: HoneySat example use cases for further extensibility.
Extensibility Use Case Relevant HoneySat Module(s) High-level Steps
Add new sensor Satellite Simulator To add a new sensor you can leverage the existing Sensor Abstract Base Class (ABC) included in the Satellite Simulator source code and write the logic for your new sensor.
Add new payload Satellite Simulator To add a new payload you can leverage the existing PayloadSimulation Abstract Base Class (ABC) included in the Satellite Simulator source code and write the logic for your new payload including where to get the payload data.
Integrate new flight software Satellite Simulator, ground configuration, satellite personality, radio simulator, flight software services To integrate a new flight software into HoneySat the first step is to analyze and understand the target flight software architecture and source code even before attempting integration. Then the target flight software must be integrated using HoneySat’s flight software services module. After integrating the flight software services, the Satellite Simulator can be connected using the Interface Python ABC locanted in the Satellite Simulator source code.
Add a new exposed network protocol Ground Software, Docker compose A new exposed network protocol can be added by modifying the Docker compose file included with HoneySat. The compose file allows for new services and TCP ports to be easily added.